OpenStudy (anonymous):

I'm trying to find "a" using the law of cosines when I have A (50 degrees), b= 13, and c = 6. The equation that they've prompted me to use is a^2 = b^2 + c^2 -2(b)(c) (cosA). The final answer I keep getting is somewhere around 26-27, but the available choices are not within that range. What am I doing incorrectly?

OpenStudy (amistre64):

is your calculator mode in degrees or radians?

OpenStudy (anonymous):

Degree.

OpenStudy (amistre64):

your calculations are in error then, im gettig something closer to 10

OpenStudy (anonymous):

Is there a certain order I would go about solving this then?

OpenStudy (amistre64):

not really ... show me yuor work

OpenStudy (anonymous):

Alright, hold on. I need to rewrite it so it's legible.

OpenStudy (amistre64):

typing is legible :)

OpenStudy (anonymous):

Alright @amistre64 sorry... Here's a link to the work. http://prntscr.com/6xat6q

OpenStudy (amistre64):

can you explain your line 3?

OpenStudy (amistre64):

yeah, your line 3 is just making stuff up ... you stopped doing math

OpenStudy (amistre64):

13^2 + 6^2 -2(13)(6) K is NOT [13^2 + 6^2 -2(13)(6)] K

OpenStudy (anonymous):

+ Line three was created through combining like terms. That is, 169 + 36 - 26.

OpenStudy (amistre64):

no it wasnt ...

OpenStudy (amistre64):

it was fabricated using imagination instead of proper math.

OpenStudy (anonymous):

If you insist.

OpenStudy (amistre64):

when does: a + bx = (a+b)x ?

OpenStudy (anonymous):

I see.

OpenStudy (amistre64):

do: 13^2 + 6^2 then subtract 2(13)(6) cos(50) then take the sqrt

OpenStudy (anonymous):

Would the 6 be by itself?

OpenStudy (amistre64):

no, 2(13)(6) cos(50) is all one term, it is a product of 4 factors.

OpenStudy (amistre64):

the order in which we multiply it out is immaterial 13 cos(50) (2)(6) or whatever

OpenStudy (amistre64):

13^2 = 169 6^2 = 36 2(13)(6) cos(50) = 156 cos(50) = 100.2748....

OpenStudy (amistre64):

169 + 36 - 100.2748... = 104.7252.... c = sqrt(104.7252....) ... which is close to 10

OpenStudy (anonymous):

I see. This makes a lot more sense, thank you.

OpenStudy (amistre64):

a = sqrt(104...) typoed the letter :) good luck